OpenStudy (anonymous):

The function f is given by F(x)=X^4+X^2-2. On which intervals is f increasing? i know im supposed to set f'x=0 but im not getting any of same answer choices.

OpenStudy (anonymous):

f'(x)=4x^3+2x=0

OpenStudy (anonymous):

Solve and factor ...

OpenStudy (anonymous):

i factored first and made it 2x(x^2+1)=0

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

solutions are x=0 and ...

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

can we solve (x^2+1)=0 or is there only one solution?

OpenStudy (anonymous):

isnt that x=1?

OpenStudy (anonymous):

Let's check 1^2 + 1 = 0??

OpenStudy (anonymous):

oh right .lol

OpenStudy (mertsj):

Find the first derivative. If it is positive, the function is increasing.

OpenStudy (anonymous):

So we have 0 as a solution .... How do we find increasing or decreasing?

OpenStudy (anonymous):

dont u plug in values above and less?

myininaya (myininaya):

factored wrong

OpenStudy (mertsj):